OpenStudy (anonymous):

Let f(x) = x − 3 and g(x) = x + 11. Find f(x) ⋅ g(x).

OpenStudy (solomonzelman):

\(\large\color{blue}{ \displaystyle f(x)=x-3 }\) \(\large\color{teal}{ \displaystyle f(x)=x+11 }\) \(\large\color{black}{ \displaystyle \color{blue}{f(x)} \times\color{teal}{g(x)}= \color{blue}{(x-3)} \times\color{teal}{(x+11)}={~~}\bf ?? }\)